# A Hybrid Multi-Factor Network with Dynamic Sequence Modeling for Early Warning of Intraoperative Hypotension

### v2

@@ 4.2 Experimental Results
Table 2 presents a comprehensive evaluation of predictive models for dynamic IOH prediction, highlighting the advantages of the HMF model. The results indicate that HMF consistently outperforms baseline methods across datasets and sampling rates. Its ability to capture the intricate dynamics of blood pressure trends is reflected in improved predictive accuracy and classification performance. Notably, HMF demonstrates robustness in handling the non-stationary and complex nature of intraoperative blood pressure data, where traditional models often struggle. In the IOH prediction task, HMF shows strong performance in detecting hypotensive events. Its improvements across different sampling rates suggest adaptability to varying temporal granularities, a crucial factor in clinical applications. Compared to LSTM and Transformer, which perform well in some scenarios but face challenges with long sequence dependencies, HMF effectively models temporal dependencies and coupling effects. Informer and DLinear, despite their advantages in general forecasting tasks, show limitations in handling intraoperative data, particularly in recall performance. Overall, HMF maintains stable performance across diverse conditions, demonstrating its potential to enhance IOH prediction in complex clinical settings.
@@
Performance Analysis w.r.t Transfer Settings.
Assessing model transferability across patients and demographic groups is essential for IOH prediction. Transfer learning enhances the robustness and generalizability of the HMF framework. To evaluate this, we apply the HMF model, trained on CH-OPBP, to new patient cohorts and age groups, comparing performance with and without transfer learning. Table 4 presents cross-patient transfer results, where a model trained on one patient is applied to another. AUC scores suggest that transfer learning preserves or even enhances predictive performance, indicating the model’s ability to capture shared physiological patterns. This reduces reliance on patient-specific data while maintaining accuracy, though individual training may still offer advantages in highly distinct cases. Table 5 presents cross-group transfer results between age cohorts. The non-transfer model performs better in elderly patients, suggesting that age-specific models better capture physiological complexities. In contrast, younger patients benefit from transfer learning, likely due to more homogeneous physiological responses. These findings underscore the importance of demographic-specific considerations while demonstrating transfer learning’s potential to improve predictive performance across diverse clinical settings.
@@
To optimize key model parameters and enhance the Patch Encoder’s ability to capture temporal dependencies, we conduct a parameter sensitivity analysis on the CH-OPBP dataset, focusing on the model dimension $d_{\text{model}}$ and patch length $S$. We first investigate the effect of varying $d_{\text{model}}$ while keeping the context window fixed at 300 and the predicted scope length at 100. As shown in Figure 5, different values of $d_{\text{model}}$ significantly influence model performance, with an optimal setting maximizing AUC. Subsequently, we examine the impact of varying $S$ while holding $d_{\text{model}}$ constant. This analysis is critical, as $S$ directly affects how effectively the model captures local temporal structures within physiological signals. The corresponding results, presented in Figure 6, indicate that selecting an appropriate $S$ value plays a crucial role in model performance. Based on these analyses, we determine the optimal parameter values that yield the highest AUC scores, ensuring robust model performance. The sensitivity of the model’s performance to $d_{\text{model}}$ and $S$ highlights the importance of appropriately configuring the Patch Encoder to accurately model blood pressure dynamics.
Visualization Case Study Analysis.
Figure 7 compares MAP predictions across models, demonstrating the effectiveness of the HMF framework. Our analysis of the CH-OPBP dataset, using a 15-minute context window and a 3-second sampling rate, consistently shows that HMF outperforms other models in predicting IOH events across various prediction horizons. Specifically, for a prediction length of 100, HMF (pink) closely follows actual MAP trends, capturing both rapid fluctuations and long-term variations. In contrast, DLinear (purple) and Informer (teal) tend to smooth out critical transitions, while Transformer (yellow) shows occasional delays. LSTM (orange) struggles with long-range dependencies, leading to deviations in extended forecasts. These results highlight HMF’s advantage in modeling complex temporal patterns for accurate IOH prediction.
Figure 7: Comparison of MAP prediction across different models on representative test cases.
@@ 5 Conclusion
We proposed a novel approach to IOH prediction by reformulating it as a time series forecasting task. To address key challenges, we introduced symmetric normalization to mitigate non-stationarity in MAP and SBP series, sequence decomposition to capture complex temporal patterns, and a patch-based Transformer model to enhance representation learning. Extensive experiments on real-world datasets validated the effectiveness of our HMF framework, demonstrating its ability to improve early IOH detection. This study highlights the potential of time series forecasting in clinical decision support and provides a foundation for future research in intraoperative risk prediction.
Appendix
Appendix A Related Work
A.1 Intraoperative Hypotension Forecasting
Existing methods for intraoperative hypotension (IOH) prediction predominantly adopt a static modeling approach, extracting handcrafted features from physiological signals rather than directly modeling their temporal dynamics. Early studies focused on high-fidelity arterial pressure waveforms, leading to the Hypotension Prediction Index (HPI) Hatib et al. [2018b]. Machine learning approaches, including ensemble methods Cherifa et al. [2020] and gradient boosting Kendale et al. [2018], integrated preoperative and intraoperative factors but largely treated data as independent samples, overlooking temporal dependencies. Deep learning models, such as RNN-based predictors Jeong et al. [2019] and attention-based frameworks Lu et al. [2023], improved performance by capturing sequential patterns but primarily relied on fixed-length feature representations, failing to fully model non-stationary hemodynamic changes. Recent interpretable models Ritter et al. [2023] enhanced clinical applicability but remained feature-driven. Our work addresses these limitations by formulating IOH prediction as a multivariate time-series forecasting problem, explicitly capturing temporal dependencies and non-stationary dynamics in arterial blood pressure signals.
A.2 Time Series and Sequence Forecasting
Time series forecasting is fundamental in domains such as finance, healthcare, and energy Cheng et al. [2024b]. Traditional models like ARIMA Ariyo et al. [2014] and exponential smoothing struggle with high-dimensional physiological signals He et al. [2023], while deep learning methods, including LSTMs Graves and Graves [2012] and GRUs Dey and Salem [2017], improve long-term dependency modeling. Recent Transformer-based models enhance long-sequence forecasting efficiency. Informer Zhou et al. [2021] optimizes self-attention mechanisms, Autoformer Chen et al. [2021] incorporates trend-seasonality decomposition, and FEDformer Zhou et al. [2022] leverages frequency domain analysis. Lightweight alternatives, such as DLinear Zeng et al. [2023] and TSMixer Ekambaram et al. [2023], demonstrate competitive performance by preserving temporal structures with simpler architectures. GPHT Liu et al. [2024] further introduces generative pretraining to improve forecasting transferability across datasets and horizons. Despite these advances Cheng et al. [2024b]; Zhang et al. [2024]; Cheng et al. [2024a]; Wang et al. [2024b, a], challenges remain in computational efficiency and real-time adaptability for physiological signal forecasting. Our work addresses these gaps by integrating multi-factor feature extraction with an efficient Transformer-based approach, improving IOH prediction accuracy in dynamic clinical settings.
Appendix B Evaluation Details
B.1 Ground Truth Labeling for IOH Events
In this study, ground truth labels for intraoperative hypotension (IOH) events were generated based on recorded mean arterial pressure (MAP) values during surgery. The labeling process identifies clinically significant hypotensive episodes while minimizing transient fluctuations that may not reflect sustained hemodynamic instability.
An IOH event is defined as a segment where the MAP value remains below a predefined threshold $\theta_{\text{MAP}}$ for at least $t$ minutes. Following clinical guidelines Wesselink et al. [2018], we set $\theta_{\text{MAP}}=65$ mmHg, a widely accepted hypotension threshold, and $t=1$ minute to ensure clinical relevance.
To generate labels, the algorithm scans the MAP time series for each patient, evaluating non-overlapping segments of length $t$. For each segment, the maximum MAP value is determined. If this value is below $\theta_{\text{MAP}}$, the entire segment is labeled as an IOH event. This approach ensures that only sustained hypotensive episodes are captured, reducing false positives from transient MAP drops. The labeling process is outlined in Algorithm 2.
Algorithm 2 Ground Truth Labeling for IOH Events
Input: $MAP_{\text{actual}}$: recorded MAP series, $t$: minimum duration for an IOH event, $T_{\text{seq}}$: total length of the MAP series.
Output: $L_{\text{actual}}$: binary label sequence for IOH events.
1:  Initialize $L_{\text{actual}}[:]=0$
2:  for $i$ in range(0, $T_{\text{seq}}-t+1$do
3:     $MAP_{\text{max}}=\max(MAP_{\text{actual}}[i:i+t])$
4:     if $MAP_{\text{max}}\leq\theta_{\text{MAP}}$ then
5:        $L_{\text{actual}}[i:i+t]=1$
6:     end if
7:  end for
8:  return $L_{\text{actual}}$
B.2 Evaluation of Hypotensive Events
To evaluate IOH detection, we compare the model’s predicted MAP trajectory against actual MAP values, assessing both event identification and duration. A predicted IOH event is correct if the forecasted MAP remains below $\theta_{\text{MAP}}$ for at least $t$ minutes.
Performance is quantified using the following metrics, including precision, recall, and F1-score:
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (13)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (14)
[…]
where TP (True Positives) are correctly identified IOH events, FP (False Positives) are incorrect predictions, and FN (False Negatives) are missed hypotensive episodes.
Additionally, we compute the time-to-detection (TTD), which measures how early the model predicts an IOH event before its actual onset. A lower TTD indicates better early-warning capability, which is essential for timely clinical intervention.
This evaluation framework ensures that our method provides meaningful predictions, distinguishing sustained hypotensive episodes from transient MAP fluctuations while maintaining high predictive accuracy.
Appendix C Details of Baseline Methods
To comprehensively evaluate the performance of HMF in early IOH prediction, we benchmarked it against a set of traditional and deep learning-based methods. These baselines were selected to cover a range of modeling approaches, from statistical methods to advanced sequence learning techniques.
C.1 Traditional Methods
We included two widely used traditional time series forecasting and classification models:
  • ARIMA Ariyo et al. [2014]: The Autoregressive Integrated Moving Average (ARIMA) model was used as a classical statistical forecasting baseline. Due to its computational intensity, we trained ARIMA on only 0.5% of the test set, balancing efficiency and representativeness in the comparison.
  • Logistic Regression: A standard binary classification model, logistic regression was applied to predict IOH events. To enhance predictive performance, time series features were extracted using the tsfresh library [^5] https://github.com/blue-yonder/tsfresh, which generated 1,566 handcrafted features, capturing statistical and frequency-domain properties.
C.2 Deep Learning Methods
We evaluated several deep learning-based baselines, each known for its effectiveness in modeling sequential dependencies in time series data:
  • LSTM Graves and Graves [2012]: The Long Short-Term Memory (LSTM) network, a recurrent neural network (RNN) variant, was included due to its capability to capture long-range temporal dependencies. Its gating mechanism mitigates vanishing gradient issues, making it suitable for sequential modeling tasks like IOH prediction.
  • GRU Dey and Salem [2017]: The Gated Recurrent Unit (GRU) is an RNN variant similar to LSTM but with a simplified architecture. GRU was included for comparison due to its efficiency and comparable performance in capturing time series patterns while using fewer parameters.
  • Transformer Vaswani et al. [2017]: The Transformer model, with its self-attention mechanism, was included to evaluate the effectiveness of attention-based architectures in IOH prediction. By allowing the model to weigh different time steps adaptively, it can capture complex dependencies beyond what RNN-based models achieve.
  • Informer Zhou et al. [2021]: The Informer model extends the Transformer architecture by introducing a ProbSparse self-attention mechanism, reducing computational cost while preserving long-range dependency modeling. Its efficiency makes it well-suited for long-sequence forecasting tasks.
  • DLinear Zeng et al. [2023]: A lightweight linear model for time series forecasting, DLinear models trend and seasonal components separately. Despite its simplicity, it has demonstrated competitive performance against complex deep learning models, providing a strong baseline.
  • TSMixer Ekambaram et al. [2023]: A recent MLP-based model, TSMixer improves time series forecasting by using a structured mixing approach for temporal dependencies. We included it to assess the performance of purely feedforward architectures compared to sequential and attention-based models.
These baselines provide a diverse comparison set, ranging from traditional statistical methods to state-of-the-art deep learning architectures. By including both simple and complex models, we ensure a rigorous evaluation of HMF’s effectiveness in IOH prediction.
Table 6: Performance Comparison of Different Methods Using The CH-OPBP Dataset.
Model Sample(s) Pred MSE MAE AUC Accuracy (%) Recall (%)
Arima 1 300 97.1580 7.5501 0.6071 90.34 23.33
600 118.0238 8.8539 0.6022 75.15 25.53
900 175.6289 10.1539 0.5795 66.48 23.13
3 100 26.9210 4.0537 0.6625 86.60 38.46
200 155.1008 10.0347 0.5870 77.27 22.22
300 156.7626 10.3933 0.5288 63.03 16.28
LR 1 300 0.5000 86.06 0.00
600 0.6777 62.13 81.90
900 0.3386 35.96 23.39
3 100 0.5977 79.84 32.50
200 0.7002 72.00 65.18
300 0.7342 75.01 65.79
LSTM 1 300 98.2363 7.9263 0.5660 90.65 14.57
600 122.3699 9.4033 0.5000 73.49 0.00
900 133.7677 9.9659 0.5226 64.47 5.18
3 100 104.2937 8.9187 0.5000 90.51 0.00
200 127.8037 10.0756 0.5000 73.48 0.00
300 141.1666 10.6498 0.5000 62.94 0.00
Transformer 1 300 106.0975 8.2478 0.6194 89.99 27.32
600 126.0585 9.6861 0.5804 75.94 19.94
900 148.2355 10.2089 0.5758 67.32 20.04
3 100 93.5870 7.4800 0.6328 89.51 30.91
200 112.6674 9.0762 0.5406 75.06 9.34
300 108.3092 8.4761 0.6177 70.22 29.11
Informer 1 300 102.4836 8.2318 0.5663 90.41 14.95
600 92.3734 7.6706 0.6952 79.20 48.91
900 116.2513 8.4248 0.6741 73.93 42.27
3 100 104.9414 8.1057 0.6065 90.11 24.28
200 108.3698 8.7226 0.5934 77.23 21.25
300 119.7905 8.3365 0.6835 74.24 45.59
DLinear 1 300 107.2025 8.2388 0.5468 90.39 10.61
600 127.4571 9.6158 0.5316 74.75 7.20
900 142.3762 10.1150 0.5208 64.33 4.82
3 100 106.1303 8.2176 0.5587 90.27 13.42
200 123.7473 9.4676 0.5711 76.31 16.21
300 141.7650 10.2002 0.5271 64.72 6.30
HMF 1 300 73.7196 6.8555 0.7702 76.48 77.70
600 89.8372 7.6161 0.7230 75.30 65.91
900 116.2464 8.2751 0.7125 74.08 60.33
3 100 77.0390 6.8708 0.7801 77.72 78.38
200 88.6844 7.4293 0.7245 75.32 66.34
300 93.7548 7.5482 0.7194 73.55 65.68
Table 7: Performance Comparison of Different Methods Using The VitalDB Dataset.
Model Sample(s) Pred MSE MAE AUC Accuracy(%) Recall(%)
Arima 3 100 193.3889 10.7640 0.5397 83.10 10.42
200 268.3629 13.7636 0.5073 60.27 5.48
300 310.3584 14.8105 0.5280 53.20 9.68
LR 3 100 0.5094 78.52 2.11
200 0.5877 61.80 37.50
300 0.5814 58.41 60.80
LSTM 3 100 157.8527 10.6642 0.5000 84.35 0.00
200 204.0663 12.3266 0.5000 61.35 0.00
300 204.2178 12.5930 0.5000 48.58 0.00
Transformer 3 100 162.8149 10.6928 0.5022 84.36 0.54
200 162.6087 10.8778 0.5006 61.40 0.13
300 150.6858 10.4997 0.5091 49.50 2.12
Informer 3 100 164.9233 11.0281 0.5000 84.35 0.00
200 152.0211 10.7310 0.5000 61.35 0.00
300 159.4176 10.9371 0.5008 48.66 0.16
DLinear 3 100 173.4594 11.1419 0.5097 84.33 2.41
200 175.6332 11.4898 0.5077 61.85 1.97
300 176.2505 11.8588 0.5049 49.08 1.19
HMF 3 100 149.5450 8.7662 0.6819 78.29 53.50
200 165.2017 9.3708 0.6370 68.01 44.74
300 182.5259 10.0166 0.6215 61.50 39.38
Appendix D Appendix D: Full Results Report
Tables 6 and 7 present the full experimental results comparing the performance of different baseline methods and our proposed HMF model on the CH-OPBP and VitalDB datasets. These tables provide a detailed evaluation across multiple metrics, including mean squared error (MSE), mean absolute error (MAE), area under the curve (AUC), accuracy, and recall, under different prediction horizons (e.g., 100, 200, and 300 sample steps). The results highlight the effectiveness of HMF in capturing intraoperative hypotension (IOH) dynamics compared to traditional and deep learning-based approaches. Notably, HMF demonstrates superior performance in key metrics such as AUC and recall, indicating its ability to better identify hypotensive events while maintaining predictive accuracy.

### v4

@@ 5.2 Experimental Results
Table 2 presents a comprehensive evaluation of predictive models for dynamic IOH prediction, highlighting the advantages of the HMF model. The results indicate that HMF consistently outperforms baseline methods across datasets and sampling rates. Its ability to capture the intricate dynamics of blood pressure trends is reflected in improved predictive accuracy and classification performance. Notably, HMF demonstrates robustness in handling the non-stationary and complex nature of intraoperative blood pressure data, where traditional models often struggle. In the IOH prediction task, HMF shows strong performance in detecting hypotensive events. Its improvements across different sampling rates suggest adaptability to varying temporal granularities, a crucial factor in clinical applications. Compared to LSTM and Transformer, which perform well in some scenarios but face challenges with long sequence dependencies, HMF effectively models temporal dependencies and coupling effects. Informer and DLinear, despite their advantages in general forecasting tasks, show limitations in handling intraoperative data. Overall, HMF maintains stable performance across diverse conditions, demonstrating its potential to enhance IOH prediction in complex clinical settings.
@@
Performance Analysis under Transfer Learning Settings.
Assessing model transferability across patients and demographic groups is essential for IOH prediction. Transfer learning enhances the robustness and generalizability of the HMF framework. To evaluate this, we apply the HMF model, trained on CH-OPBP, to new patient cohorts and age groups, comparing performance with and without transfer learning. Table 4 presents cross-patient transfer results, where a model trained on one patient is applied to another. AUC scores suggest that transfer learning preserves or enhances predictive performance, indicating the model’s ability to capture shared physiological patterns. This reduces reliance on patient-specific data while maintaining accuracy, though individual training may still offer advantages in distinct cases. Table 5 presents cross-group results between age cohorts. The non-transfer model performs better in elderly patients, suggesting that model better capture physiological complexities. In contrast, younger patients benefit from transfer learning, likely due to more homogeneous physiological responses. These findings underscore the importance of demographic considerations while demonstrating transfer learning’s potential to improve performance across diverse clinical settings.
@@
To optimize key parameters and enhance the Patch Encoder’s ability to capture temporal dependencies, we perform sensitivity analysis on the CH-OPBP dataset, focusing on model dimension $d_{\text{model}}$ and patch length $S$. We first vary $d_{\text{model}}$ with a fixed context window of 300 and prediction length of 100. As shown in Figure 5, performance is sensitive to $d_{\text{model}}$, with a clear optimum maximizing AUC. We then examine the effect of varying $S$ while keeping $d_{\text{model}}$ constant. Since $S$ governs the model’s ability to capture local temporal patterns, its selection is important. Results in Figure 6 confirm that appropriate $S$ values significantly affect forecasting accuracy. These findings guide the selection of optimal hyperparameters to ensure performance and highlight the importance of configuring the Patch Encoder to effectively model blood pressure dynamics.
@@ 6 Conclusion
In this paper, we proposed a novel approach to intraoperative hypotension (IOH) prediction by reformulating it as a dynamic sequence forecasting task, enabling proactive and continuous risk assessment. To address key challenges inherent in modeling arterial blood pressure dynamics, we introduced symmetric normalization to mitigate non-stationarity in mean arterial pressure (MAP) and systolic blood pressure (SBP) series, sequence decomposition to better capture complex temporal patterns, and a patch-based Transformer model to enhance representation learning and forecasting accuracy. Extensive experiments on real-world intraoperative datasets validated the effectiveness and robustness of our HMF framework, demonstrating its strong performance in improving early IOH forecasting under various clinical conditions.
Acknowledgements
This research was supported by grants from the grants of Provincial Natural Science Foundation of Anhui Province (No.2408085QF193), USTC Research Funds of the Double First-Class Initiative (No. YD2150002501), the Fundamental Research Funds for the Central Universities of China(Grant No. PA2024GDSK0112, No. WK2150110032), Anhui Provincial Natural Science Foundation (No. 2308085MG226), the Key Technologies R&D Program of Anhui Province (No. 202423k09020039).